# SPECIAL RELATIVITY

### v9

@@ 2.5 Myth about the Constancy of the Speed of Light
It is generally believed that experiments show that the speed of light in vacuum is independent of the source or observer (4). This statement presented in most textbooks and is obviously incorrect. The constancy of the speed of light is related to the choice of synchronization convention, and cannot be subject to experimental tests (5).
@@
In fact, in order to measure the one-way speed of light one has first to synchronize the infinity of clocks assumed attached to every position in space, which allows us to perform time measurements. Obviously, an unavoidable deadlock appears if one synchronizes the clocks by assuming a-priori that the one-way speed of light is $c$. In fact, in that case, the one-way speed of light measured with these clocks (that is the Einstein speed of light) cannot be anything else but $c$: this is because the clocks have been set assuming that particular one-way speed in advance.
@@
Therefore, it can be said that the value of the one-way speed of light is just a matter of convention without physical meaning. In contrast to this, the two-way speed of light, directly measurable along a round-trip, has physical meaning, because round-trip experiments rely upon the observation of simultaneity or non-simultaneity of events at a single point in space and not depends on clock synchronization convention. All well known methods to measure the speed of light are, indeed, round-trip measurements. The cardinal example is given by the Michelson-Morley experiment: this experiment uses, indeed, an interferometer where light beams are compared in a two-way fashion.
@@
Because of our usage of Galilean transformations within electrodynamics we have some apparent paradoxes. One finds many books which say that a Galilean transformations of the velocity of light is not consistent with the electron-theoretical explanation of refraction and reflection (6). It is widely accepted that if we consider a moving source and a stationary glass, the incident light wave and wave scattered by the dipoles of the glass cannot interfere as required by the electron theory of dispersion since their velocity are different. This is misconception. It is clear that an incident wave with a certain frequency, no matter what its velocity, excites the electrons of a glass into oscillations of the same frequency. They then emit radiation with the same frequency. Thus, the incident and scattered wave at any given point have the same frequency and can interfere. The effect of the different velocities is to produce a relative phase which varies with position in space.
@@ 2.6 Convention-Dependent and Convention-Invariant Parts of the Theory
Consider the motion of charged particle in a given magnetic field. The theory of relativity says that the particle trajectory $\vec{x}(t)$ in the lab frame depends on the choice of a convention, namely the synchronization convention of clocks in the lab frame. Whenever we have a theory containing an arbitrary convention, we should examine what parts of the theory depend on the choice of that convention and what parts do not. We may call the former convention-dependent, and the latter convention-invariant parts. Clearly, physically meaningful measurement results must be convention-invariant.
@@
Consider the motion of two charged particles in a given magnetic field, which is used to control the particle trajectories. Suppose there are two apertures at point $A$ and at point $A^{\prime}$. From the solution of the dynamics equation of motion we may conclude that the first particle gets through the aperture at $A$ and the second particle gets through the aperture at $A^{\prime}$ simultaneity. The two events, i.e. the passage of particles at point $A$ and point $A^{\prime}$ have exact objective meaning i.e. convention-invariant. However, the simultaneity of these two events is convention-dependent and has no exact objective meaning. It is important at this point to emphasize that, consistently with the conventionality of simultaneity, also the value of the speed of particle is a matter of convention and has no definite objective meaning.
@@
In order to examine what parts of the dynamics theory depend on the choice of that convention and what parts do not, we want to show the difference between the notions of path and trajectory. Let us consider the motion of a particle in three-dimensional space using the vector-valued function $\vec{x}(t)$. We have a prescribed curve (path) along which the particle moves. The motion along the path is described by $l(t)$, where $l$ is a certain parameter (in our case of interest the length of the arc). The trajectory of a particle conveys more information about its motion because every position is described additionally by the corresponding time instant. The path is rather a purely geometrical notion. If we take the origin of the (Cartesian) coordinate system and we connect the point to the point laying on the path and describing the motion of the particle, then the creating vector will be a position vector $\vec{x}(l)$ (7).
The difference between trajectory $\vec{x}(t)$ and path $\vec{x}(l)$ is very interesting. The path has exact objective meaning i.e. it is convention-invariant. In contrast to this, and consistently with the conventionality intrinsic in the velocity, the trajectory $\vec{x}(t)$ of the particle is convention dependent and has no exact objective meaning.
@@
In order to avoid being to abstract for to long we have given some examples: just think of the experiments related with accelerator physics. Suppose we want to perform a particle momentum measurement. A uniform magnetic field can be used in making a ”momentum analyzer” for high-energy charge particles, and it must be recognized that this method for determining the particle’s momentum is convention-independent. In fact, the curvature radius of the path in the magnetic field (and consequently the three-momentum) has obviously an objective meaning, i.e. is convention-invariant. Dynamics theory contains a particle trajectory that we do not need to check directly, but which is used in the analysis of electrodynamics problem.
@@ 2.7 Myth about the Reality of Relativistic Time Dilation and Length Contraction
Generally, experts on the theory of relativity erroneously identify the properties of Minkovski space-time with the familiar form that certain convention-dependent quantities assume under the standard Lorentz coordinatization. These quantities usually are called ”relativistic kinematics effects”. There is a widespread belief that the convention-dependent quantities like the time dilation, length contraction, and Einstein’s addition of velocities have direct physical meaning. We found that statement like ” moving clocks run slow” is not true under the adopted absolute time clock synchronization, and, hence, are by no means intrinsic features of Minkowski space-time. Relativistic kinematic effects are coordinate (i.e. convention-dependent) effects and have no exact objective meaning (8). In the case of Lorentz coordinatization, one will experience e.g. the time dilation phenomenon. In contrast to this, in the case of absolute time coordinatization there are no relativistic kinematics effects and no time dilation will be found. However, all coordinate-independent quantities like the particle path $\vec{x}(l)$ and momentum $|\vec{p}|$ remain independent of such a change in clock synchronization.
@@ 2.8 Relativistic Particle Dynamics
2.8.1 Covariant Approach
The accelerated motion is described by a covariant equation of motion for a relativistic charged particle under the action of the four-force in the Lorentz lab frame. The trajectory of a particle $\vec{x}_{cov}(t)$ is viewed from the Lorentz lab frame as a result of successive infinitesimal Lorentz transformations. The lab frame time $t$ in the equation of motion cannot be independent from the space variables. This is because Lorentz transformations lead to a mixture of positions and time, and the relativistic kinematics effects are considered to be a manifestation of the relativity of simultaneity.
2.8.2 Noncovariant (”Single Frame”) Approach
Let us consider the conventional particle tracking approach. It is generally accepted that in order to describe dynamics of relativistic particles in the lab reference frame, which we assume inertial, one only needs to take into account the relativistic dependence of the particle momentum on the velocity. The treatment of relativistic particle dynamics involves only corrected Newton’s second law. In a given lab frame, there is an electric field $\vec{E}$ and magnetic field $\vec{B}$. They push on a particle in accordance with
@@ 2.8.2 Noncovariant (”Single Frame”) Approach
This solution of the dynamics problem in the lab frame makes no reference to Lorentz transformations. Conventional particle tracking treats the space-time continuum in a non-relativistic format, as a (3+1) manifold. In other words, in this approach, introducing as only modification to the classical case the relativistic mass, time differ from space. In fact, we have no mixture of positions and time (9).
@@
Most of the interesting phenomena in which charges move under the action of electromagnetic fields occur in very complicated situations. But here we just want to discuss the simple problem of the accelerated motion of particles in a constant magnetic field. According to the non-covariant treatment, the magnetic field is only capable of altering the direction of motion, but not the speed (i.e. mass) of a particle. This study of relativistic particle motion in a constant magnetic field, usual for accelerator engineering, looks precisely the same as in nonrelativistic Newtonian dynamics and kinematics. The trajectory of a particle $\vec{x}(t)$, which follows from the solution of the corrected Newton’s second law, does not include relativistic kinematics effects as relativity of simultaneity and the Galilean vectorial law of addition of velocities is actually used.
@@
Let us discuss the important problem of the addition of velocities in relativity. Suppose that in the case of accelerated motion one introduces an infinite sequence of co-moving frames. At each instant, the rest frame is a Lorentz frame centered on the particle and moving with it. Suppose that in inertial frame where particle is at rest at a given time, the traveler was observing light itself. In other words measured speed of light $v=c$, and yet the frame is moving relative the lab frame. How will it look to the observer in the lab frame? According to relativistic law of addition of velocities the answer will be $c$. Maxwell’s equations remain in the same form when Lorentz transformations are applied to them, but Lorentz transformations give rise to non-Galilean transformation rules for velocities, and therefore the theory of relativity shows that, if Maxwell’s equations is to be valid in the lab frame, the trajectories of the particles must include relativistic kinematics effects. In other words, Maxwell,s equations can be applied in the lab frame only in the case when particle trajectories are viewed, from the lab frame as the result of successive infinitesimal Lorentz transformations.
@@
The absence of relativistic kinematics effects is the prediction of conventional non-covariant theory and is obviously absurd from the viewpoint of Maxwell’s electrodynamics. Therefore, something is fundamentally, powerfully, and absolutely wrong in coupling fields and particles within a ”single inertial frame”.
@@ 2.9 Mistake in Commonly Used Method of Coupling Fields and Particles
It is generally believed that the electrodynamics problem can be treated within the same ”single inertial frame” description without reference to Lorentz transformations. In all standard derivations it is assumed that usual Maxwell’s equations and corrected Newton’s second law can explain all experiments that are performed in a single inertial frame, for instance the lab reference frame.
Going to electrodynamics problem, the differential form of Maxwell’s equations describing electromagnetic phenomena in the same inertial lab frame (in cgs units) is given by the following expressions:
@@
where $\delta(\vec{x}-\vec{x}_{n}(t))$ is three-dimensional delta function, $m_{n},e_{n},\vec{x}_{n}(t)$, and $\vec{v}_{n}=d\vec{x}_{n}(t)/dt$ denote the mass, charge, position, and the velocity of the $n$th particle, respectively. To evaluate radiation fields arising from an external sources Eq. (8) we need the velocity $\vec{v}_{n}$ and the position $\vec{x}_{n}$ as a function of lab frame time $t$. It is generally accepted by physics community that equation of motion, which describes how coordinates of the particle carrying the charge change with time $t$, is described by corrected Newton’s second law Eq. (4).
@@
This coupling of Maxwell’s equations and corrected Newton’s equation is commonly accepted as useful method in accelerator physics and, in particular, in analytical and numerical calculations of radiation properties. Such approach to relativistic dynamics and electrodynamics usually forces the accelerator physicist to believe that the design of particle accelerators possible without detailed knowledge of the theory of relativity.

### v10

@@ 2.5 Myth about the Constancy of the Speed of Light
It is generally believed that experiments show that the speed of light in a vacuum is independent of the source or observer [^4] Many authors of textbooks still attribute a measurable status to the conventional quantities. To quote e.g. Cristodoulides [7] ”The fact that Galilean transformation does not leave Maxwell’s equations has already been mentioned […] On the other hand, experiments show that the speed of light in a vacuum is independent of the source or observer.”. This statement is presented in most textbooks and is incorrect. The constancy of the speed of light is related to the choice of synchronization convention, and cannot be subject to experimental tests  [^5] Because we have empirical access only to the round-trip average speed of light, statements about the magnitude and isotropy of the one-way speed of light must reflect the assumptions made in the choice of time coordinatization, and such entities change as the theory is re-synchronized. To quote C. Anderson, I. Vetharaniam, and G. Stedman [8]: ”No experiment, then, is a ”one-way” experiment. An empirical test of any property of the one-way speed of light is not possible. Such quantities as the one-way speed of light are irreducibly conventional in nature, and recognizing this aspect is to recognize a profound feature of nature”..
@@
To measure the one-way speed of light one has first to synchronize the infinity of clocks assumed attached to every position in space, which allows us to perform time measurements. Obviously, an unavoidable deadlock appears if one synchronizes the clocks by assuming a-priori that the one-way speed of light is $c$. In fact, in that case, the one-way speed of light measured with these clocks (that is the Einstein speed of light) cannot be anything else but $c$: this is because the clocks have been set assuming that particular one-way speed in advance.
@@
Therefore, it can be said that the value of the one-way speed of light is just a matter of convention without physical meaning. In contrast to this, the two-way speed of light, directly measurable along a round-trip, has physical meaning, because round-trip experiments rely upon the observation of simultaneity or non-simultaneity of events at a single point in space and not depends on clock synchronization convention. All well-known methods to measure the speed of light are, indeed, round-trip measurements. The cardinal example is given by the Michelson-Morley experiment: this experiment uses, indeed, an interferometer where light beams are compared in a two-way fashion.
@@
Because of our usage of Galilean transformations within electrodynamics, we have some apparent paradoxes. One finds many books that say that Galilean transformations of the velocity of light are not consistent with the electron-theoretical explanation of refraction and reflection  [^6] The peculiarity of the kinematic consequence of using Galilean transformations is that the speed of light emitted by a moving source depends on the relative velocity between the source and the observer. A widespread theoretical argument used to support the incorrectness of Galilean transformations is the conclusion that a Galilean transformation of the velocity of light is not consistent with the explanation of reflection and refraction. This idea is a part of the material in well-known books. To quote Pauli [9] ”[…] it is essential that the spherical waves emitted by the dipoles in the body should interfere with the incident wave. If we now think of the body as at rest, and the light source moving relative to it, then […] the wave emitted by the dipoles will have a velocity different from that of the incident wave. Interference is therefore not possible.”. It is widely accepted that if we consider a moving source and a stationary glass, the incident light wave and wave scattered by the dipoles of the glass cannot interfere as required by the electron theory of dispersion since their velocity are different. This is a misconception. It is clear that an incident wave with a certain frequency, no matter what its velocity, excites the electrons of a glass into oscillations of the same frequency. They then emit radiation with the same frequency. Thus, the incident and scattered wave at any given point have the same frequency and can interfere. The effect of the different velocities is to produce a relative phase that varies with position in space.
@@ 2.6 Convention-Dependent and Convention-Invariant Parts of the Theory
Consider the motion of a charged particle in a given magnetic field. The theory of relativity says that the particle trajectory $\vec{x}(t)$ in the lab frame depends on the choice of a convention, namely the synchronization convention of clocks in the lab frame. Whenever we have a theory containing an arbitrary convention, we should examine what parts of the theory depend on the choice of that convention and what parts do not. We may call the former convention-dependent and the latter convention-invariant parts. Clearly, physically meaningful measurement results must be convention-invariant.
@@
Consider the motion of two charged particles in a given magnetic field, which is used to control the particle trajectories. Suppose there are two apertures at point $A$ and at point $A^{\prime}$. From the solution of the dynamics equation of motion, we may conclude that the first particle gets through the aperture at $A$ and the second particle gets through the aperture at $A^{\prime}$ simultaneity. The two events, i.e. the passage of particles at point $A$ and point $A^{\prime}$ have exact objective meaning i.e. convention-invariant. However, the simultaneity of these two events is convention-dependent and has no exact objective meaning. It is important at this point to emphasize that, consistently with the conventionality of simultaneity, also the value of the speed of the particle is a matter of convention and has no definite objective meaning.
@@
To examine what parts of the dynamics theory depend on the choice of that convention and what parts do not, we want to show the difference between the notions of path and trajectory. Let us consider the motion of a particle in three-dimensional space using the vector-valued function $\vec{x}(t)$. We have a prescribed curve (path) along which the particle moves. The motion along the path is described by $l(t)$, where $l$ is a certain parameter (in our case of interest the length of the arc). If we take the origin of the (Cartesian) coordinate system and connect the point to the point laying on the path, then the creating vector will be a position vector $\vec{x}(l)$  [^7] For a general discussion of the difference between path and trajectory we suggest reading the book [10].. The difference between trajectory $\vec{x}(t)$ and path $\vec{x}(l)$ is very interesting. The path has exact objective meaning i.e. it is convention-invariant. In contrast to this, and consistently with the conventionality intrinsic in the velocity, the trajectory $\vec{x}(t)$ of the particle is convention-dependent and has no exact objective meaning.
@@
Let us give some examples: just think of the experiments related to accelerator physics. Suppose we want to perform a particle momentum measurement. A uniform magnetic field can be used in making a ”momentum analyzer” for high-energy charged particles, and it must be recognized that this method for determining the particle’s momentum is convention-independent. The curvature radius of the path in the magnetic field (and consequently the three-momentum) has an objective meaning, i.e. is convention-invariant. Dynamics theory contains a particle trajectory that we do not need to check directly, but which is used in the analysis of electrodynamics problems.
@@ 2.7 Myth about the Reality of Relativistic Time Dilation and Length Contraction
Generally, experts on the theory of relativity erroneously identify the properties of Minkowski space-time with the familiar form that certain convention-dependent quantities assume under the standard Lorentz coordinatization. These quantities usually are called ”relativistic kinematics effects”. There is a widespread belief that convention-dependent quantities like time dilation, length contraction, and Einstein’s addition of velocities have direct physical meaning. Relativistic kinematic effects are coordinate (i.e. convention-dependent) effects and have no exact objective meaning  [^8] The standard textbooks erroneously identify the properties of space-time with the familiar form that certain synchronization-dependent quantities assume under Einstein’s clock synchronization. These quantities are, for example, the formulae for ”time dilation” and ”length contraction”. Fortunately, it has also been occasionally stressed in the literature that the forms of the above ”relativistic effects” are coordinate-dependent, their forms depend in turn on the kind of synchronization procedure adopted. To quote Leubner, K. Auflinger, and P. Krumm [11]: ”… there is a widespread belief among students that the familiar form of coordinate-dependent quantities like the measured velocity of light, the Lorentz transformation between two observers, ”addition of velocities”, ”time dilation”, ”length contraction” which they assume under the standard clock synchronization, is relativity proper. To demonstrate that this is by no means so, this paper studies the consequences of a non-standard synchronization, and it is shown that drastic changes in the appearance of all these quantities are thus induced.” . In the case of Lorentz coordinatization, one will experience e.g. the time dilation phenomenon. In contrast to this, in the case of absolute time coordinatization, there are no relativistic kinematics effects, and no time dilation will be found. However, all coordinate-independent quantities like the particle path $\vec{x}(l)$ and momentum $|\vec{p}|$ remain independent of such a change in clock synchronization.
@@ 2.8 Relativistic Particle Dynamics
The accelerated motion is described by a covariant equation of motion for a relativistic charged particle under the action of the four-force in the Lorentz lab frame. The trajectory of a particle $\vec{x}_{cov}(t)$ is viewed from the Lorentz lab frame as a result of successive infinitesimal Lorentz transformations. The lab frame time $t$ in the equation of motion cannot be independent of the space variables. This is because Lorentz transformations lead to a mixture of positions and time, and the relativistic kinematics effects are considered to be a manifestation of the relativity of simultaneity.
Let us consider the conventional particle tracking approach. It is generally accepted that in order to describe the dynamics of relativistic particles in the lab reference frame, which we assume inertial, one only needs to take into account the relativistic dependence of the particle momentum on the velocity. The treatment of relativistic particle dynamics involves only corrected Newton’s second law. In a given lab frame, there is an electric field $\vec{E}$ and magnetic field $\vec{B}$. They push on a particle in accordance with
@@
This solution of the dynamics problem in the lab frame makes no reference to Lorentz transformations. Conventional particle tracking treats the space-time continuum in a non-relativistic format, as a (3+1) manifold. In other words, in this approach, introducing as an only modification to the classical case the relativistic mass, the time differs from space. We have no mixture of positions and time  [^9] Clarification of the true content of the non-covariant theory can be found in various advanced textbooks. To quote e.g. Ferrarese and Bini [12]: ”… within a single inertial frame, the time is an absolute quantity in special relativity also. As a consequence, if no more than one frame is involved, one would not expect differences between classical and relativistic kinematics. But in the relativistic context, there are differences in the transformation laws of the various relative quantities (of kinematics or dynamics), when passing from one reference frame to another.” We see that authors give a special role to the concept of a ”single inertial frame”. The name ”single inertial frame” tends to suggest that a distinctive trait of non-covariant theory is the absence of relativistic kinematics in the description of particle motion. This point was expressed by Friedman [15]: ”Within any single inertial frame, things look precisely the same as in Newtonian kinematics: there is an enduring Euclidean three-space, a global time $t$, and the law of motion.”.
@@
Most of the interesting phenomena in which charges move under the action of electromagnetic fields occur in very complicated situations. But here we just want to discuss the simple problem of the accelerated motion of particles in a constant magnetic field. According to the non-covariant treatment, the magnetic field is only capable of altering the direction of motion, but not the speed (i.e. mass) of a particle. This study of relativistic particle motion in a constant magnetic field, usual for accelerator engineering, looks precisely the same as in non-relativistic Newtonian dynamics and kinematics. The trajectory of a particle $\vec{x}(t)$, which follows from the solution of the corrected Newton’s second law, does not include relativistic kinematics effects as relativity of simultaneity and the Galilean vectorial law of addition of velocities is used.
@@
Let us discuss the important problem of the addition of velocities in relativity. It is clear that the light velocity transforms as a particle velocity. We cannot take one kinematics for a particle’s velocity and the other kinematics for the velocity of light. We consider a case when the light source in the lab frame is accelerated from rest up to velocity $u$. Suppose that in the Lorentz inertial frame where the light source is at rest, the traveler was observing light itself. In other words, $v=c$, and yet the frame is moving. How will it look to the observer in the Lorentz lab frame? According to the relativistic law of addition of velocities, the answer will be $v=c$. Maxwell’s equations remain in the same form when Lorentz transformations are applied to them, but Lorentz transformations give rise to non-Galilean transformation rules for velocities, and therefore the theory of relativity shows that, if Maxwell’s equations are to be valid in the lab frame, the trajectories of the particles must include relativistic kinematics effects. In other words, Maxwell’s equations can be applied in the lab frame only in the case when particle trajectories are viewed, from the lab frame as the result of successive infinitesimal Lorentz transformations.
@@
The absence of relativistic kinematics effects is the prediction of conventional (3+1) non-covariant particle tracking and is obviously absurd from the viewpoint of Maxwell’s electrodynamics. The relativity of simultaneity (i.e. mixture of positions and time) does not have a place in this description. Therefore, something is fundamentally, powerfully, and absolutely wrong in coupling fields and particles within a ”single inertial frame”.
@@ 2.9 Mistake in Commonly Used Method of Coupling Fields and Particles
It is generally believed that the electrodynamics problem can be treated within the same ”single inertial frame” description without reference to Lorentz transformations. In all standard derivations, it is assumed that Maxwell’s equations and corrected Newton’s second law can explain all experiments that are performed in a single inertial frame, for instance, the lab reference frame.
Going to the electrodynamics problem, the differential form of Maxwell’s equations describing electromagnetic phenomena in the same inertial lab frame (in cgs units) is given by the following expressions:
@@
where $\delta(\vec{x}-\vec{x}_{n}(t))$ is three-dimensional delta function, $m_{n},e_{n},\vec{x}_{n}(t)$, and $\vec{v}_{n}=d\vec{x}_{n}(t)/dt$ denote the mass, charge, position, and the velocity of the $n$th particle, respectively. To evaluate radiation fields arising from external sources Eq. (8) we need the velocity $\vec{v}_{n}$ and the position $\vec{x}_{n}$ as a function of lab frame time $t$. It is generally accepted by the physics community that the equation of motion, which describes how coordinates of the particle carrying the charge change with time $t$, is described by corrected Newton’s second law Eq. (4).
@@
This coupling of Maxwell’s equations and corrected Newton’s equation is commonly accepted as a useful method in accelerator physics and, in particular, in analytical and numerical calculations of radiation properties. Such an approach to relativistic dynamics and electrodynamics usually forces the accelerator physicist to believe that the design of particle accelerators is possible without detailed knowledge of the theory of relativity.